two people, 7220 ft. apart, observe a balloonist flying overhead between them. their measures of the angles of elevation at 3:30pm are 35.6° and 58.2° respectively. what is the height of the balloonist above the ground?
1
Expert's answer
2019-05-27T12:59:51-0400

Let we have triangle ABC, where AC is the distance between two people (7220 ft), AB is a distance between one of them and baloonist, and angle BAC is equal to 35.6 degrees. BC is a distance from another person to the baloonist, and angle BCA is equal to 58.2 degrees. Let point H be a projection of baloonist to the AC, and BH is a height of the balloonist above the ground.


(1)"tan(BAH)=BH\/AH\u27f9BH=tan(BAH)\u2217AH",

(2)"tan(BCH)=BH\/(AC\u2212AH)\u27f9BH=tan(BCH)\u2217(AC\u2212AH)"


from these two equations we have


"tan(BAH)\u2217AH=tan(BCH)\u2217(AC\u2212AH)", so


"AH=AC*tan(BCH)\/(tan(BAH)+tan(BCH))"


let use it for equation (1)


"BH=AC*tan(BAH)*tan(BCH)\/(tan(BAH)+tan(BCH))"


"BH=7220(ft)*tan(35.6 degrees)*tan(58.2 degrees)\/(tan(35.6 degrees)+tan(58.2 degrees))"


"BH=7220(ft)*0.7159*1.6128\/(0.7159+1.6128)"


"BH=3579.8(ft)"